OpenStudy (anonymous):

How about 3x+7=-3x+2+6x ?

Nnesha (nnesha):

combine like terms on right side 6c - 3x = ???

OpenStudy (anonymous):

-9x+2 on the right side and 3x + 7 on the left side right?

Nnesha (nnesha):

nope 6 - 3 is what ???

OpenStudy (anonymous):

\[3 x+7=6 x-3 x+2 \]\[3 x+7=3 x+2 \]\[7=2 \]An absurdity.